# On Subjective Uncertainty Quantification and Calibration in Natural Language Generation

### v1

@@ 2.4 Representing and Eliciting Epistemic Uncertainty
We now turn to the problem of epistemic uncertainty quantification in in-context learning, where the prompt $I=\langle x_{1},y_{1},\ldots,x_{n},y_{n},x_{*}\rangle$ is formed by $n$ demonstrations $(x_{i},y_{i})$ followed by a test input $x_{*}$. We assume that $r(y,y^{\prime};I)=r(y,y^{\prime};x_{*})$ only depends on $I$ through $x_{*}$, which is commonly satisfied in applications. We propose the following measure of epistemic uncertainty inspired by Fong et al., (2024); Xu and Raginsky, (2022):
$\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*})$
$\hskip 40.00006pt-\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*}).$ (7)
In the above, $p_{x}$ denotes a user-specified distribution of additional ICL inputs, $y_{n+1:N}$ are model generations, $N>n$ is a hyperparameter, and we assume the LM has good performance for ICL up to a sample size of $N$. Eq. (7) can be estimated through autoregressive sampling. As long as we use the same set of samples for its both terms, the resulted estimate will always be non-negative, due to the convexity of the functional $\rho\mapsto\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{\rho(y_{*})}r(y^{\prime},y_{*};x_{*})$.
@@
To understand (7), first consider a simplified scenario as follows: suppose the LM $p_{M}$ is equivalent to a Bayesian model that assumes $\{(x_{i},y_{i})\}$ are i.i.d. conditional on a latent variable $\theta$, $p_{x}$ defines i.i.d. $x_{n+1:N}$, and $\{x_{n+i},y_{n+i}\}$ are sufficiently informative so that the likelihood function $p(y=\cdot\mid x=x_{*},\theta)$ determined by $\theta\mid x_{1:N},y_{1:N}$ is identifiable given infinite samples. Then under mild technical conditions, as $N\to\infty$ (7) will become equivalent to
$\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*})-\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*}).$ (7’)
@@
The equivalence is a consequence of Doob, (1949)’s theorem and can be proved with the same idea as Fong et al., (2024). For completeness, we provide a proof in Appendix A. (7) is a subjective notion of excess risk. In other words, it is the amount of reducible risk should we have full knowledge about the distribution of $y\mid x=x_{*}$, which can be determined given either the latent $\theta$, or an infinite amount of missing data $\{x_{>n},y_{>n}\}$. As noted in Xu and Raginsky, (2022), this is precisely the intuition behind Bayesian epistemic uncertainty, and we recover standard epistemic uncertainty measures (e.g., posterior variance for a regression mean, or mutual information) if we return to simple scenarios where $r$ is the square or log loss.
@@
Within the Bayesian framework, (7) with a finite choice of $N$ will provide a lower bound for (7): by convexity the second term may become larger. The intuition is that knowledge of the samples $\{x_{n+1:N},y_{n+1:N}\}$ does not always allow for the full reduction of risk. It is possible to view the lower bound as an indication of the “true epistemic uncertainty” (7), and it can be tightened if we choose $\{x_{n+i}\}$ to be more similar to $x_{*}$. The bound will also be tight regardless of the choice of $N$ if the uncertainty is “fully aleatoric” or “fully epistemic”, i.e., if the risk for $y\mid x=x_{*}$ cannot be reduced by any number of additional observations $(x_{n+1:N},y_{n+1:N})$ or can be fully reduced using a single $(x_{n+1},y_{n+1})$. Alternatively, we can observe that we do not always need the tightest bound for (7): if all we can do is to collect $N-n$ real samples, the remaining proportion of the risk will be effectively irreducible.
@@
As discussed in §1, the Bayesian view above may be justified through a common belief that high-capacity LMs may approximate Bayesian inference. It is also supported by the results of Wen et al., (2022, §2.6, §4), which imply that in a multi-task setting, any $p_{M}$ with a strong average-case performance for prediction will yield a similar value for (7) as the true Bayesian model. Falck et al., (2024) investigated this belief empirically on synthetic low-dimensional datasets, and found it to be valid for smaller choices of $N$, although it breaks down as $N\to\infty$ becomes larger.333 Falck et al., (2024) also studied epistemic uncertainty quantification for ICL but restricted to datasets where the method of Fong et al., (2024) is applicable. Their method assumes $\{x_{n+1:N},$ $y_{n+1:N}\}$ can be modelled by a low-dimensional statistical model, so that a posterior for the model parameter can be constructed and visualized. This is unrealistic for free-form natural language data. Taken together, these results suggest that the Bayesian perspective could be relevant when we restrict $N$ to a scope where ICL is sample efficient.
@@
Independent to the Bayesian perspective, Eq. (7) can also be understood through its connection to recent works. Collier et al., (2022) studied a broadly similar problem where a proportion of the combined uncertainty can be reduced by conditioning on privilege information; this leads to a similar measure for reducible uncertainty as (7), if we view $(x_{n+1:N},y_{n+1:N})$ as the latent “privilege information”. Johnson et al., (2024); Ahdritz et al., (2024) proposed methods that measure the correlation between consecutive predictions $(y_{n+1},y_{*})$ given the same input $x_{n+1}=x_{*}$; as argued in Johnson et al., (2024), any learning system $p_{M}$ with complete confidence for the true data distribution should lead to conditionally i.i.d. $(y_{*},y_{n+1})$ given $I$, and any $p_{M}$ with a “fully epistemic” uncertainty should produce $y_{*}=y_{n+1}$. It is clear that (7) matches these behaviours at the extremes: it equals 0 in the former case, and correctly indicate a full reduction of risk in the latter case (i.e., it equals $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ assuming $r(y,y;I)\equiv 0$).The Bayesian perspective provides additional insights to the works of Johnson et al., (2024); Ahdritz et al., (2024). Eq. (7) generalizes their methodology by allowing for a wider range of choices for $\{x_{n+1:N}\}$, and the above discussion provides guidance on their choices.
Finally, we note that while (7) is generally applicable to MBR decoding, with a trivial choice of $\mathcal{Y}_{I}^{\prime}$ it may take a trivial value: if we use the Gibbs predictor which always samples from $p_{M}(\cdot\mid x_{1:n},y_{1:n})$, we will find (7) to be zero; this is to be expected as the predictor does not make any attempt to utilize the additional information. In our experiments we choose $\mathcal{Y}_{I}^{\prime}$ to contain multiple samples, one from each sample path for $y_{n+1:N}$. As we will see, this is sufficient to provide meaningful estimates of epistemic uncertainty.
@@ 3 Experiments
3.1 Free-Form Question Answering
We first illustrate the proposed method on two QA tasks chosen by Lin et al., (2024): CoQA (Reddy et al.,, 2019) and NQOpen (Lee et al.,, 2019). CoQA contains reading comprehension questions where the answer can generally be deduced from the prompt, whereas NQOpen is an open-domain task and requires information not present in the prompt. Kalai and Vempala, (2024) proved in simplified scenarios that instruction-tuned LMs may have to be uncalibrated on the latter type of task; we investigate their theoretical predictions empirically.
Figure 1: Reliability diagrams for GPT-3.5 on QA tasks. Error bar denotes 95% confidence interval for the average observed utility.
We follow the setup in Lin et al., (2024) and adopt the GPT-3.5 model (OepnAI,, 2024), both for generation and to define the automated evaluation criterion; the latter determines our utility function and in turn the uncertainty measure (2).444 In this aspect our method differs from Lin et al., (2024), where uncertainty measures are defined with estimated entailment relations or lexical metrics. The evaluation criterion is defined using ICL, by instructing the LM to rate the consistency between two responses based on few-shot demonstrations (Lin et al.,, 2024, App. B.3). We reuse the model samples released by Lin et al., (2024), which contains 20 sampled responses for each query. Following their work we focus on the Gibbs predictor.
The results are shown in Fig. 1. As we can see, the LM’s subjective uncertainty generally correlates well with the average observed utility. It is slightly overconfident on the reading comprehension task, but much more so on the open-domain task. Such a difference appears consistent with the theoretical analysis of Kalai and Vempala, (2024).
3.2 In-Context Machine Translation
We now consider machine translation with ICL, where utility is defined through the widely adopted chrF score (Popović,, 2015). chrF measures both semantic and syntactic differences; the latter can be relevant in translation due to the need to detect unnatural or invalid syntactic structures. The inherent syntactic differences in the possible generations for $y\mid x$ constitute a source of irreducible uncertainty, making machine translation a non-trivial scenario to illustrate our uncertainty decomposition method.
We follow the setup of Agarwal et al., (2024) who demonstrated that ICL can lead to steady improvements for machine translation. We adopt the FLORES+ dataset (NLLB Team et al.,, 2022), in particular the translation tasks from English to the following languages: French, Tamil, Yue Chinese (Cantonese), and Tigrinya. We use the following LMs: GPT-3.5, Gemini-1.5-Flash, and Gemini-1.5-Pro (Gemini Team,, 2024), and vary the number of ICL shots $n\in\{4,16\}$. For each task, we subsample 128 queries $(x_{*},y_{*})$ from the datasets. For each query, we draw 35 samples from the LM to approximate its predictive distribution, and another 5 samples to define the action space $\mathcal{Y}^{\prime}_{I}$. For the quantification of epistemic uncertainty, we set $N=n+8$, and sample $x_{n+1:N}$ by instructing an LM to rewrite the test query $x_{*}$; see Appendix B for details.
$n$ gemini-1.5-flash gemini-1.5-pro gpt-3.5
$4$ 0.276 / 0.356 0.332 / 0.431 0.185 / 0.329
$16$ 0.246 / 0.371 0.304 / 0.426 0.168 / 0.351
Table 1: Machine translation (English $\to$ Yue Chinese): gECE (Eq. (6), $\downarrow$) and chrF ($\uparrow$) for varying choices of $n$.
We first evaluate the calibration of the combined subjective uncertainty on the Yue Chinese task. As shown in Table 1, the LMs are generally poorly calibrated, although calibration improves as the sample size $n$ increases. Calibration plots are presented in Figure 3 in Appendix, which shows that all LMs are overconfident. It is interesting to note an anticorrelation between the task performance and calibration among the three LMs evaluated. We note that Yue Chinese may constitute a low-resource translation task (Haddow et al.,, 2022) for which the LMs’ pretraining corpus is likely less informative. As is also shown in Figure 3, calibration is better on the French task for which more corpus is available.
Figure 2: Machine translation (English $\to$ Yue Chinese): reduction of epistemic uncertainty (EU) vs total uncertainty (TU) as the ICL sample size increases from $4$ to $16$.
We now investigate whether the LMs possess a coherent notion of epistemic and aleatoric uncertainty, in particular whether epistemic uncertainty is the only reducible component as sample size increases. For this purpose we plot the reduction of total uncertainty (TU, (2)) against that of epistemic uncertainty (EU, (7)) on each data point. Figure 2 shows that the two quantities have a good match. Deviation from perfect equivalence is expected here, due to the variance in the Monte-Carlo estimation for EU and TU.
Language French Tamil Yue Tigrinya
Avg. EU 0.016 0.059 0.078 0.091
chrF 0.737 0.392 0.329 0.084
# speakers (M) 312 87 87 10
Table 2: Machine translation: average epistemic uncertainty on different languages, obtained from gpt-3.5 using $n=4$ ICL shots. We also provide the chrF score and the number of speakers (from Ethnologue,, 2024, in million) of each language as an indication of its linguistic resource available.
Finally, we evaluate if the epistemic uncertainty measure (7) is meaningful, by comparing its average value across different translation tasks: we expect sensible notions of epistemic uncertainty to be higher on tasks for low-resource languages, as they are likely less represented in the pretraining corpus. Table 2 confirms this intuition: the LM demonstrates higher epistemic uncertainty on lower-resource languages.

### v2

@@ 2.4 Representing and Eliciting Epistemic Uncertainty
We now turn to the quantification of epistemic uncertainty, using ICL as a key example. Epistemic uncertainty is the proportion of uncertainty that can be reduced given additional observations. It is particularly relevant when such observations are costly to obtain or incorporate into the generation process. In these scenarios, we can limit their use to queries with a higher level of reducible (i.e., epistemic) uncertainty.
Epistemic uncertainty in ICL.
In ICL the LM is prompted with $I=\langle z_{1:n},x_{*}\rangle=\langle x_{1},y_{1},\ldots,x_{n},y_{n},x_{*}\rangle$, where $z_{i}:=(x_{i},y_{i})$ are demonstration input-output pairs and $x_{*}$ denotes the test query. $n$ can be arbitrary. We assume that the user’s utility does not depend on $z_{1:n}$ given $x_{*}$, so that the risk has the form $r(y,y^{\prime};I)\equiv r(y,y^{\prime};x_{*})$. We propose the following measure of epistemic uncertainty, inspired by Fong et al., (2023); Xu and Raginsky, (2022):
$\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{x}(x_{n+1:N}\mid I)}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\mathbb{E}_{p_{M}(y_{*}\mid z_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*})$
$\hskip 40.00006pt-\mathbb{E}_{p_{x}(x_{n+1:N}\mid I)}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(y_{*}\mid z_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*}).$ (5)
In the above, $p_{x}$ denotes a user-specified distribution of inputs, $y_{n+1:N}\sim\prod_{j=n+1}^{N}p_{M}(y_{j}\mid z_{1:j-1},x_{j})$ follow the LM’s predictive distribution, and $N>n$ is a hyperparameter. (5) can be estimated through autoregressive sampling and Monte-Carlo estimation; see §4.2 for implementation details.
Comparing with (2), we can see that (5) is the difference between the total uncertainty from two predictors; the second predictor has access to the additional demonstrations $z_{n+1:N}=(x_{n+1:N},y_{n+1:N})$, where $y_{n+1:N}$ are model generations. Thus, (5) is a subjective measure of the reducible uncertainty given $x_{n+1:N}$. By convexity of the $\inf$ functional, (5) is always non-negative; so is its Monte-Carlo estimate as long as we use the same set of samples for its both terms.
Bayesian justifications.
@@
To further understand (5), consider a simplified setting as follows: suppose $p_{M}$ is equivalent to a Bayesian model that assumes $z_{1:n}$ are i.i.d. conditional on a latent variable $\theta$, $x_{n+1:N}\sim p_{x}(\cdot\mid I)$ are conditionally i.i.d., and $z_{n+1:N}$ are sufficiently informative so that the likelihood function $p(y=\cdot\mid x=x_{*},\theta)$ determined by $\theta\mid z_{1:N}$ is identifiable given infinite samples. Then under mild technical conditions, as $N\to\infty$ (5) will become equivalent to
$\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(\theta\mid z_{1:n})}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*})-\mathbb{E}_{p_{M}(\theta\mid z_{1:n})}\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*}).$ (5’)
@@
The equivalence can be proved with the same idea as Fong et al., (2023); we provide a proof in App. A for completeness. (5) represents the amount of reducible risk should we have full knowledge about the distribution of $y\mid x=x_{*}$, which can be determined given either the latent $\theta$ or an infinite amount of missing data. As noted in Xu and Raginsky, (2022), This is precisely the intuition behind Bayesian epistemic uncertainty, and we recover standard epistemic uncertainty measures (e.g., posterior variance for a regression mean, or mutual information) if we return to the standard choices of $r$ such as the square or log loss.
@@
Within the Bayesian framework, (5) with a finite choice of $N$ will provide a lower bound for (5): by convexity the second term will be larger. The intuition is that knowledge of the samples $z_{n+1:N}$ does not always enable the full reduction of risk. We can view the lower bound as an indication of the “true” epistemic uncertainty (5), and it can be tightened if we choose $\{x_{n+i}\}$ to be more similar to $x_{*}$. The bound will also be tight for finite $N$ if the uncertainty is “fully aleatoric” or “fully epistemic”, i.e., if the risk at $x=x_{*}$ cannot be reduced by any amount of additional $z_{n+1:N}$ or if it can be fully reduced by a single $z_{n+1}$. Alternatively, we can observe that we do not always need the tightest bound for (5): if all we can do is to collect $N-n$ real samples, the remaining uncertainty will be effectively irreducible.
@@
As discussed in §1, this Bayesian view may be justified through a common belief that high-capacity LMs may approximate Bayesian inference. It is also supported by the results of Wen et al., (2022, §2.6, §4), which imply that in a multi-task setting, any $p_{M}$ with a strong average-case performance for prediction will yield a similar value for (5). Falck et al., (2024) investigated this belief empirically on synthetic numerical datasets and found it to be valid for smaller choices of $N$. Taken together, these results suggest that the Bayesian perspective can be relevant when we restrict $N$ to a regime where ICL is statistically efficient.
Connection to non-Bayesian methods.
@@
Independent to the above, Eq. (5) can also be understood through its connection to recent works. Johnson et al., (2024); Ahdritz et al., (2024) proposed methods that measure the correlation between consecutive predictions $(y_{n+1},y_{*})$ for the same input $x_{n+1}=x_{*}$; as argued in Johnson et al., (2024), any learning system $p_{M}$ with complete confidence should lead to conditionally i.i.d. $(y_{*},y_{n+1})$ given $I$, and any $p_{M}$ with a “fully epistemic” uncertainty should produce $y_{*}=y_{n+1}$. It is clear that (5) matches these behaviours at the extremes: it equals 0 in the former case, and in the latter case correctly indicate a fully reducible risk. The Bayesian perspective provides additional insights to the works of Johnson et al., (2024); Ahdritz et al., (2024). Eq. (5) generalizes their methodology by allowing for a wider range of choices for $\{x_{n+1:N}\}$, and the above discussion provides guidance on their choices.
Remark 2.2 (generalizations).
(5) can be readily applied beyond ICL if we use $(x_{n+1:N},y_{n+1:N})$ to represent more general queries, e.g., any clarifying questions and their answers. In such scenarios, (5) remains a non-negative measure of reducible uncertainty. While we will focus on ICL in experiments, we note that this generalization can be useful when the real responses are more expensive to obtain compared to LM generations, for example when they require interaction with user (Li et al.,, 2023) or a third-party source (Min et al.,, 2024).
3 Related Work
Probabilistic UQ for LMs.
UQ has long been studied in LM applications involving a short or fixed-form response (e.g., Kamath et al.,, 2020; Jiang et al.,, 2021). Free-form generation introduces additional challenges. Our discussions are motivated by recent works on semantic uncertainty (Kuhn et al.,, 2023; Tian et al.,, 2023), which we seek to clarify and extend to a general scope. The measure (2) is a generalized entropy (DeGroot,, 1962); similar to Xu et al., (2019, p.8), we deviate from that line of classical work by considering general loss functions and computational constraints.
All methods reviewed so far can be viewed as adopting a probabilistic perspective: they quantify uncertainty through the variation in the LM’s predictive distribution. Another line of work proposes to prompt the LM for its uncertainty, which we discuss shortly. Orthogonal to these developments, conformal prediction methods (e.g., Quach et al.,, 2023; Mohri and Hashimoto,, 2024; Yadkori et al., 2024b, ) take a prespecified uncertainty measure and a calibration set as input and provide a predictor with coverage guarantees.
Calibration in free-form generation.
Band et al., (2024) studied calibration in scenarios where the free-form generation $y$ is used for downstream classification based on a known predictor $\phi(y)$. Our work generalizes their setup, as we can always define a utility for $y$ based on $\phi$ and the standard utility functions for classification. Huang et al., 2024a studied the rank-calibration of general uncertainty measures which, as discussed in their work, is a different property. We note that an LM could be arbitrarily overconfident or underconfident for the uncertainty measure (2) to remain rank-calibrated. Huang et al., 2024b proposed several calibration metrics based on a similarity measure. In their work both the similarity measure and the method to aggregate calibration error are left underspecified, rendering the resulted metrics hard to interpret; in contrast, our Eq. (4) admits a straightforward interpretation as a generalized ECE.
Prompt-based methods, instruction-tuned LMs.
A recent line of work proposes to quantify uncertainty through prompting, e.g., by asking the LM to output the probability that its generation is correct (Xiao and Wang,, 2021; Mielke et al.,, 2022; Ren et al.,, 2023) or to answer a multiple-choice question (Kadavath et al.,, 2022). Compared with the probabilistic methods, this approach can be more efficient computationally as it avoids the need to draw multiple samples for each query. However, it requires the LM to have stronger reasoning capabilities, and is more difficult to apply when the user’s utility involves factors beyond factual correctness or for the quantification of epistemic uncertainty.
Still, within the above scope, it is reasonable to ask whether the verbalized approach might be preferable for instruction-tuned LMs. The question remains largely open: previous studies have found that instruction tuning can affect probabilistic calibration (Achiam et al.,, 2023; Tian et al.,, 2023), but empirical comparisons between probabilistic and verbalized approaches yielded inconsistent results (Kuhn et al.,, 2023; Tian et al.,, 2023; Xiong et al.,, 2024). Moreover, in the free-form setting there is also a lack of principled evaluation of calibration: many prior works (Kuhn et al.,, 2023; Lin et al.,, 2024) adopted evaluation metrics that conflate calibration with predictive performance (Huang et al., 2024a, , p.5), and some studies (Jiang et al.,, 2021; Tian et al.,, 2023) evaluated token-level (rather than task-specific) calibration, which is less relevant. Our experiments in §4.1 will fill in this gap.
Epistemic uncertainty.
Our method is inspired by Fong et al., (2023); Xu and Raginsky, (2022). Similar ideas were discussed by Berti et al., (2021); Wen et al., (2022) in relation to Fong et al., (2023), and Grünwald and Dawid, (2004); Kotelevskii and Panov, (2024) in relation to Xu and Raginsky, (2022). Similar notions of reducible uncertainty also appear in active learning (Smith et al.,, 2023), and in experiment design (Rainforth et al.,, 2024) in relation to Remark 2.2. All of these works either restricted to the estimation of a full data distribution, or required explicitly Bayesian models,333 i.e., a $p_{M}$ that corresponds to a Bayesian posterior predictive, defined by a known posterior over model parameters and a likelihood. rendering them inapplicable to general LM applications. Falck et al., (2024) applied the method of Fong et al., (2023) to ICL, but similarly restricted to density estimation for numerical data and approximately Bayesian models. Concurrent work of Yadkori et al., 2024a proposed a measure similar to (5), but based on the log loss. It thus cannot distinguish between task-specific and irrelevant uncertainties, and also requires access to the LM’s predictive density; our work avoids such limitations.
@@ 4 Experiments
4.1 Uncertainty in Question Answering
We first illustrate the uncertainty and calibration measures by revisiting the question answering (QA) experiments in recent works (Lin et al.,, 2024; Tian et al.,, 2023). As discussed in §3, our goal is to determine whether probabilistic UQ methods are suitable for modern instruction-tuned LMs, by providing a principled evaluation of task-specific calibration.
Experiment setup.
We combine the QA datasets studied by Lin et al., (2024); Tian et al., (2023): CoQA (Reddy et al.,, 2019), TriviaQA (Joshi et al.,, 2017), NQOpen (Kwiatkowski et al.,, 2019), SciQ (Welbl et al.,, 2017), and TruthfulQA (Lin et al.,, 2021). We subsample 1000 queries for each dataset and report bootstrap confidence intervals (CIs) for all metrics. The utility $S(y^{\prime},y;I)$ is defined by instructing an LM to rate the consistency of $y^{\prime}$ with $y$ as an answer to the user’s question. We adopt the prompt templates in Lin et al., (2024) for evaluation and generation. We add a system prompt that instruct LMs to generate single-line responses, and use gpt-4o-mini for evaluation.444 We confirmed that evaluation remains aligned with human perception, and that except on SciQ the generations are still varied in form. SciQ is a short-form dataset which we include for consistency with Tian et al., (2023). As TriviaQA and NQOpen contain trivia questions that were annotated well before the knowledge cutoff date of current LMs, we further remove a subset of questions with clearly outdated answers; see App. B.1 for details.
We use LMs from the following model families: gpt-4o (OpenAI,, 2024), llama-3.1 (Dubey et al.,, 2024), gemini-1.5 (Gemini Team,, 2024), and claude-3.5 (Anthropic,, 2024). All LMs are instruction tuned unless otherwise noted. In a subset of experiments, we compare the uncertainty measure (2) (denoted as Prob.) with the prompt-based methods of Kadavath et al., (2022, denoted as P(True)) and Tian et al., (2023, Verb.). We apply all methods to the Gibbs predictor, and estimate uncertainty measures using Monte-Carlo estimation with 10 samples for each query. To evaluate calibration, we use the ECE metric (4) estimated through histogram binning (Naeini et al.,, 2015). For reference, we also compute the average test utility to quantify predictive performance, as well as the AURAC metric from Nadeem et al., (2009); Lin et al., (2024), which measures the performance of a selective prediction procedure based on a given LM and uncertainty measure. Full setup details are deferred to App. B.1.
Results and discussion.
We defer full results to App. B.2 and summarize the main findings below.
Figure 1: Question answering: ECE and utility for the llama-3.1-70b model before and after instruction tuning. Error bar denotes 95% bootstrap CI.
How does instruction tuning affect calibration?
We revisit an experiment in Tian et al., (2023) on the impact of instruction tuning on token-level calibration. Fig. 1 plots the ECE and test utility of llama-3.1-70b before and after instruction tuning. Its left subplot can be compared with Tian et al., (2023, Fig. 1 left, for the first three datasets). The comparison reveals a rather different phenomenon: whereas the calibration of token-level probability deteriorated significantly (an increase in ECE by 0.10.10.10.10.20.20.20.2) across all three datasets in Tian et al., (2023), we find the deterioration of task-specific calibration to be much smaller, except on TruthfulQA. It is interesting to note that among these datasets, TruthfulQA is the only one where instruction-tuning led to a significant improvement in utility; additional results on NQOpen and CoQA further confirm this trend. These results indicate that the impact of instruction tuning on calibration may have been overestimated in previous work. Results for AUARC are provided in App. B.2, where we also evaluate additional LMs. We find that across all datasets, the performance of selective prediction based on task-specific uncertainty is either approximately unchanged or improved after instruction tuning. This is similarly in contrast to Tian et al., (2023, Fig. 1 right), which shows that selective prediction based on token-level uncertainty consistently performs worse after instruction tuning.
Figure 2: Question answering: test utility vs ECE for isntruction-tuned LMs, averaged over the datasets in Lin et al., (2024). App. B.2 presents results for individual datasets.
We now compare the calibration of instruction-tuned LMs with their predictive performance. As shown in Fig. 2, calibration is generally correlated with predictive performance, but substantial variation remains. It is thus reasonable to expect that the calibration of many instruction-tuned LMs may still be improved with different tuning methodologies.
Figure 3: Question answering: calibration error from different methods. Error bar denotes 95% bootstrap CI.
Comparing probabilistic and prompt-based uncertainty.
We compare the calibration of the uncertainty measure (2) with prompt-based alternatives, for which we can define an analogous ECE metric. Fig. 3 visualizes ECE from different methods. We can see that the measure (2) generally leads to the best calibration error; the only exception is on TruthfulQA, which is “adversarially constructed” to evaluate LMs’ tendency to repeat human falsehoods. Reliability diagrams in App. B.2 provide further insights. Taken together, the results suggest that the probabilistic approach to UQ remains effective for modern instruction-tuned LMs.
Note that in the above comparison, only the probabilistic uncertainty measure makes use of the utility $S$. However, as discussed before, in many scenarios it is reasonable to assume access to (a surrogate of) $S$. Moreover, the ability of the probabilistic approach to incorporate different choices of utility can also be viewed as an advantage, especially when utility cannot be interpreted as a probability of correctness.
Figure 4: Machine translation: average utility vs the number of deferrals to the many-shot predictor. We also report the p-value of a permutation test that compares the AUC-DF from the EU method to random.
4.2 Epistemic Uncertainty in ICL
We now turn to epistemic uncertainty (EU) quantification in ICL, and investigate whether the proposed EU measure may provide an effective deferral strategy that routes selected queries to a more expensive predictor.
Background and setup.
LM generations can often be improved given information retrieved from an external source. Such a process can become expensive if retrieval is from a third party (Min et al.,, 2024) or if it leads to a very long prompt (Agarwal et al.,, 2024). In such scenarios, it would be desirable to limit retrieval to queries with a higher level of reducible uncertainty.
Agarwal et al., (2024) showed that many-shot ICL can be effective in low-resource machine translation (Haddow et al.,, 2022). Informed by their findings, we evaluate the use of EU for choosing between a LM predictor given few-shot demonstrations and a more expensive predictor based on many-shot ICL. We compare this method (denoted as EU) to a random baseline, and an alternative based on total uncertainty (TU, Eq. (2)).
We follow the setup in Agarwal et al., (2024) and adopt the chrF score (Popović,, 2015) as the utility. chrF is a lexical similarity score that measure both semantic and syntactic differences; the latter provides a source of irreducible uncertainty. We define a base predictor using ICL with $n=4$ shots, and a more expensive predictor with $n^{\prime}=128$. We defer to the expensive predictor on $m$ test queries with the highest EU, and report the average utility (over all test samples) w.r.t. $m$. Following Gupta et al., (2024), we refer to this function as a deferral curve, and use its integral (AUC-DF) as a scalar summary.
The EU measure (5) is instantiated with $N-n=4$ and estimated as follows: for a query $I=\langle z_{1:n},x_{*}\rangle$, we first draw $l$ i.i.d. samples $\{z^{(i)}_{n+1:N}=(x^{(i)}_{n+1:N},y^{(i)}_{n+1:N})\}_{i=1}^{l}$ where $x^{(i)}_{n+1:N}\sim p_{x}(\cdot|I)$ and $y^{(i)}_{n+1:N}\sim\prod_{k=n+1}^{N}p_{M}(y_{k}\mid\langle z_{1:n},z_{n+1:k-1}^{(i)},x_{k}^{(i)}\rangle)$. Then, for each $i$, we sample $\{y^{(i,j)}_{*}\}_{j=1}^{m}\sim p_{M}(y_{*}\mid\langle z_{1:n},z_{n+1:N}^{(i)},x_{*}\rangle)$. We use $l(m-1)$ samples to define a plug-in Monte-Carlo estimator, and the remaining samples, $\{y^{(i,m)}_{*}\}_{i=1}^{l}$, to define the action space in (5). We define $p_{x}$ by using an LM to rewrite $x_{*}$, and choose $l=5,m=8$.
Following Agarwal et al., (2024), we adopt the translation tasks from English to Kurdish (kmr) and Tamil (tam) in the FLORES+ dataset (NLLB Team et al.,, 2022). Due to resource limitations, we subsample 500 data points from each dataset. We consider the following LMs: gpt-4o-mini, gpt-4o and gemini-1.5-pro. Full setup details are deferred to Appendix B.3.
Result and discussion.
Fig. 4 plots the deferral curves from the GPT models. We can see that the EU method consistently outperforms the random baseline, and the improvement in AUC-DF is statistically significant. In contrast, the TU baseline is not always better than random, which is expected since a larger number of in-context demonstrations will not help if predictive uncertainty is irreducible. To further compare between the EU and TU methods, we compute the bootstrap distributions of their AUC-DF metrics, which are obtained by resampling the test queries without replacement. As shown in Table 1, for all but one experiment (gemini-1.5-pro on kmr) EU outperforms TU in more than 75% of bootstrap simulations.
Table 1: Machine translation: AUC-DF difference between EU and TU methods. We report the median, $25\%$ and $75\%$ percentiles across 1000 bootstrap simulations.
gemini-1.5-pro gpt-4o-mini gpt-4o
tam $0.05$ $[0.00,0.09]$ $0.09$ $[0.01,0.16]$ $0.12$ $[0.06,0.19]$
kmr $0.02$ $[-0.02,0.06]$ $0.15$ $[0.10,0.21]$ $0.15$ $[0.08,0.22]$
In the gemini-1.5-pro experiments, both EU and TU demonstrate less improvement over the random baseline. This may be attributable to the fact that uncertainty from gemini-1.5-pro is less calibrated: the respective base predictor achieves an average ECE of 0.303 on the two datasets, compared with 0.104 using gpt-4o-mini and 0.083 using gpt-4o. Nonetheless, EU still appears to be the more effective method, as indicated by Table 1. See App. B.4 for full results.